evaluate ∫ (a tan x + b cot x)^2 dx

evaluate
∫ (a tan x + b cot x)^2 dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (atanx+bcotx)^{2}dx
I = \int (a^2tan^2x+b^2cot^2x+2ab)dx
I = \int (a^2(sec^2x-1)+b^2(cosec^2x-1)+2ab)dx
I = \int a^2sec^2xdx+\int b^2cosec^2xdx-\int (a-b)^2dx
I = a^2tanx+\int b^2cosec^2xdx-\int (a-b)^2dx
I = a^2tanx- b^2cotx-\int (a-b)^2dx
I = a^2tanx- b^2cotx- (a-b)^2x+c